OpenStudy (anonymous):

the cost of eleven cds and one dvds is $69.41; and the cost of seven dvds and 12 cds is $ 112.11. How much do a cd and dvd cost?

OpenStudy (anonymous):

11C+D=$69.41 12C+7D=$112.11 where C is the cost of one CD and D is the cost of one DVD Then you just solve that system of equations for C and D. You know how to do that?

OpenStudy (lilg132):

cd = 5.75 dvd = 6.16

OpenStudy (anonymous):

\[C = 9344/1625\approx5.75015=$5.75, D = 40029/6500\approx6.15831=$616\]

OpenStudy (lilg132):

u forgot the decimal point for d's answer :p